# Optimal performance of a three stroke heat engine in the microscopic regime

### v1

@@ Abstract
In the microscopic regime, we consider a stroke-based heat engine that utilizes the temperature difference between two heat baths to produce work using a working body. The working body of the engine aims to extract heat from the hot heat bath, generate work, and discharge the surplus heat into the cold heat bath through the successive execution of three strokes. We discuss the thermodynamic framework for such an engine that involves any arbitrary-dimensional working body. Next, our investigation focuses on the smallest version of this engine, which involves a two-level working body, resulting in a closed expression for optimal work production and efficiency. Furthermore, we investigate the possibility of realizing such an engine within the Jaynes-Cummings model, gaining insights into optimal work production and efficiency. Finally, we scale down the engine, taking into account finite-dimensional heat baths, and obtain the expressions for optimal work and efficiency in this scenario.
@@ I Introduction
Within the domain of thermodynamics, heat engines play a pivotal role by converting thermal energy into work. The study of heat engines not only propels technological advancements in energy conversion but also deepens our understanding of nature’s fundamental physical laws. However, the advent of quantum mechanics has ushered in a new era of insight, particularly concerning the foundational underpinnings of thermodynamics on a microscopic scale. This captivating avenue of inquiry traces its origin back to the 1950s, when investigations into the thermodynamic properties of lasers commenced [1, 2, 3]. Since then, a resounding interest has flourished in comprehending the mechanisms of quantum thermal engines as well as quantum thermal machines in general [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21], and in formulating microscopic thermodynamic frameworks to characterize their operations [22, 23, 24, 25, 26, 27]. Recently, researchers have realized the microscopic thermal heat engines in the lab using a trapped single calcium ion as a working body [28, 29, 30, 31, 32], as well as in superconducting circuits [33], nitrogen vacancy centers in diamonds [34] atomic [35] and phononic systems [36], as well as single-electron transistors [37, 38].
@@
Microscopic heat engines can be described as thermodynamic devices that involve small-sized components. In particular, the microscopic heat engines described in the literature involve working bodies that are finite-dimensional and can be as simple as two-level systems [23, 15]. Furthermore, the analysis of the microscopic heat engines has been extended further by considering heat baths of finite size [39, 40, 41, 42, 43]. Based on their way of functioning, the heat engines considered in the microscopic regime can be broadly classified into two classes: continuous engines [11, 7, 44, 45, 46, 47] and stroke-based engines [48, 49, 50, 51]. In a continuous engine, the working body is always in contact with the hot and cold heat baths. As a consequence, the engine withdraws heat from the hot heat bath and converts it into work in a continuous and simultaneous manner. On the other hand, in a stroke-based heat engine, the working body exchanges heat and converts it into work in discrete strokes that are well separated in time. Therefore, in stroke-based engines, the working body repeatedly couples between the hot and cold baths using subsequent strokes.
The complete characterization of a continuous heat engine necessitates the consideration of various models of heat baths and couplings. These models are typically formalized through master equations, which, in turn, depend on multiple parameters such as coupling constants, bath spectral densities, etc [52, 53]. Consequently, optimizing the work extraction and efficiency of a continuous heat engine proves to be a formidable task. Despite the inherent complexity, a more accessible understanding of the basic operations of these continuous heat engines can be attained by considering stroke-based heat engines. Importantly, this approach doesn’t require detailed knowledge of the dynamics. By embracing this perspective, one can acquire significant knowledge about the fundamental limitations, optimal work, and efficiency of continuous thermal machines.
@@
In light of this, in this paper we examine a stroke-based heat engine that operates in the microscopic regime. By microscopic, we are referring to the fact that only two-level systems have been taken into account as working bodies. In the first stroke the working body aims to withdraw heat from the hot heat bath, followed by the second and third stroke where the working body converts the heat into the work and disposes the extra amount of heat into cold heat bath and completes the cycle. In order to prevent any unaccounted energy sources from contributing to the work done by the engine, we assume that the interaction between the working body and heat baths is energy-conserving. In other words, the engine generates work by harnessing the temperature difference between the two heat baths. This led us to describe the transformation of the state of the working body via a thermal operation [54, 26, 55]. The process of extracting work is represented by a unitary process, in which the working body of the engine converts a portion of the heat extracted from the hot heat bath into work, while the remaining heat is deposited to a cool heat bath. We provide the thermodynamic framework for this specific heat engine. In particular, we show that the efficiency of such an engine is always upper bounded by the Carnot efficiency. In this work, we introduce a straightforward way to optimize the efficiency of such an engine in a more generalized setting, where the range of thermal operations with respect to a heat bath is restricted. Considering the restricted range of the thermal operations, we can investigate the feasibility of implementing this engine in a more realistic scenario, such as the Jaynes-Cummings interaction. Furthermore, we extend our study to a scenario in which the heat bath’s dimension is finite.
@@ II.2.1 Heat stroke
This is the initial stroke of the engine, in which the working body interacts with the hot heat bath at thermal equilibrium using energy-saving unitary $U_{S,H}$, i.e, $[U_{S,H},H_{S}+H_{H}]=0$. Thus, the transformation of the state $\rho_{S}$ is governed by a thermal operation ${\cal E}_{H}$ [56], i.e.,
@@ II.2.2 Work stroke
In the second stroke of the engine, the working body, which resulted from the first stroke, attempts to produce work using the heat it drew from the first stroke. The engine extracts work using a unitary operation that transforms the working body’s state from the first stroke to one with lower average energy. Therefore, this stroke transforms the working body’s state as follows:
$\mathcal{E}_{H}(\rho_{S})\rightarrow V(\mathcal{E}_{H}(\rho_{S}))V^{\dagger}.$ (6)
The amount of extracted work is given by
$W=\text{Tr}\Big{(}H_{S}\big{(}\mathcal{E}_{H}(\rho_{S})-V(\mathcal{E}_{H}(\rho_{S}))V^{\dagger}\big{)}\Big{)},$ (7)
where $V$ is any arbitrary unitary. The above expression clearly shows that the unitary, which orders the eigenvalues of the density matrix $\mathcal{E}_{H}(\rho_{S})$ in decreasing order, will maximize the extracted amount. In other words, the expression of work done by the engine as given in Eq. (7) will be maximized for unitary, which leads to the passive version of the state $\mathcal{E}_{H}(\rho_{S})$ [8]. This maximum amount of the extracted work is defined as ergotropy of the state $\rho_{S}$, i.e.,
@@
$\leq$ $\max_{V\in\mathbb{U}(d)}\text{Tr}\Big{(}H_{S}\big{(}\mathcal{E}_{H}(\rho_{S})-V(\mathcal{E}_{H}(\rho_{S}))V^{\dagger}\big{)}\Big{)}$ (9)
$=$ $\text{Tr}\Big{(}H_{S}\big{(}\mathcal{E}_{H}(\rho_{S})-\mathcal{E}_{H}^{\downarrow}(\rho_{S})\big{)}\Big{)}=R(\mathcal{E}_{H}(\rho_{S})),$ (10)
where $\mathcal{E}_{H}^{\downarrow}(\rho_{S})$ denotes the passive version of the state $\mathcal{E}_{H}(\rho_{S})$, $R(\mathcal{E}_{H}(\rho_{S}))$ denotes the ergotropy of the state $\mathcal{E}_{H}(\rho_{S})$ and $\mathbb{U}(d)$ is the set of $d$-dimensional unitary [8, 57]. Next, we shall describe the final stroke of the working body.
@@ II.2.3 Closing the thermodynamic cycle
In the final stroke of the engine, the state of the working body that results from the work stroke $V{\cal E}_{H}(\rho_{S})V^{\dagger}$ interacts with the cold heat bath in thermal equilibrium via an energy conserving unitary $U_{S,C}$, i.e., $[U_{S,C},H_{S}+H_{C}]=0$. This transformation corresponds to a thermal operation with respect to the cold heat bath denoted by ${\cal E}_{C}$, i.e.,
@@ II.3 The open cycle engine
Figure 1: The above figure depicts the equivalence between the proposed three-stroke engine and open cycle engine introduced in Sec. II.3. On the left panel we have shown the proposed three stroke engine where the first stroke and the third stroke aiming for withdrawing heat from the hot bath and disposing the excess heat in the cold bath, whereas the production of work happened during the second stroke. The first and third strokes are characterized by the thermal operations at the hot and cold inverse temperatures $\beta_{H}$ and $\beta_{C}$, respectively. This proposed engine is equivalent to an open cycle engine where the third stroke is characterized by thermalization at some effective inverse temperature $\beta^{v}_{C}$ as shown in the right panel. This equivalence is capturing the underlying idea behind the proof of theorem 2.
@@ II.4 Mathematical tools
This section provides a brief overview to the mathematical tools associated with thermal operations that are important to evaluate the performance of the engine. A thermal operation can be defined as a CPTP map ${\cal E}$, which transforms a system’s state $\rho$ as
@@
where $U$ satisfies the relation $[U,H_{S}+H_{E}]=0$ with $H_{S}$, $H_{E}$ are the Hamiltonians of the system and bath, respectively, and $\gamma_{E}=\frac{e^{-\beta H_{E}}}{\text{Tr}(e^{-\beta H_{E}})}$ is the thermal state of the bath that is thermalized at an inverse temperature $\beta$.
The central task is then described by the following interconversion problem: what is the set of final states that can be achieved from a given state $\rho$ via thermal operations? This set is convex and referred to as the thermal cone of the state $\rho$, denoted by ${\cal T}(\rho)$, which follows straightforwardly from the convexity of thermal operation [58]. Characterizing the necessary and sufficient conditions for the interconversion problem in general is challenging and unknown beyond dimension two [59, 60]. Nonetheless, the seminal work by Horodecki and Oppenheim in Ref. [56] characterizes the necessary and sufficient conditions for the state interconversion problem when $[\rho,H_{S}]=0$. These states are referred to as energy-incoherent states as they do not possess coherence in the energy eigenbasis. For a given energy incoherent state $\rho$, the thermal cone ${\cal T}(\rho)$ is reduced to a polytope, i.e., the set will have a finite number of extremal points. In this case, we refer the set ${\cal T}(\rho)$ as the thermal polytope. The thermal operation that transforms the initial energy incoherent state $\rho$ to an extremal point of the thermal polytope ${\cal T}(\rho)$ is called an extremal thermal operation. Later, we shall see that these extremal points of the set ${\cal T}(\rho)$ and the extremal thermal operations play a crucial role in optimizing the performance of the three stroke engine.
@@
Here, we would like to make an important remark. When we define thermal operations, we suppose that our system of interest is in touch with a very large bath that has an effectively infinite volume. In such an ideal scenario, all possible states within the set ${\cal T}(\rho)$ can be realized. In a more practical context, various restricted subsets of states emerge within the thermal polytope ${\cal T}(\rho)$ as a result of constraints involving the feasibility of thermal operations. In such cases, the range of $\lambda$ in Eq. (21) is more generally, confined between
@@ III.2 Optimal performance of the three stroke engine
The complete proof of this theorem can be found in Appendix C which requires the concept of $\beta$-ordering that we have described in Appendix B [26, 55].
Let us analyze the scenario where one can perform any thermal operation at inverse temperature $\beta_{C}$ and $\beta_{H}$ in the first and third strokes, respectively. This implies $\lambda^{\text{max}}_{H}=1$ and $\lambda^{\text{max}}_{C}=1$. In that scenario, the optimal work produced by the engine and the optimal efficiency from Eqs. (29) and (30) are reduced to [66]
@@
respectively. Therefore, one must follow the protocol in Box 1 to achieve the best work production and efficiency of the stroke-based discrete heat engine with the hot and cold heat baths at inverse temperatures $\beta_{H}$ and $\beta_{C}$, respectively.
@@
Figure 2: Plot of efficiency vs $\beta_{C}/\beta_{H}$ with $\omega=0.2/\beta_{H}$. The solid, dashed, dotdashed, and dotted curves represent optimal efficiency, efficiencies for finite heat bath with $d=15$, $d=10$, and $d=5$, respectively.
Figure 3: Plot of $\beta_{H}$ork vs $\beta_{C}/\beta_{H}$ with $\omega=0.2/\beta_{H}$. The solid, dashed, dotdashed, and dotted curves represent optimal work production, work productions for finite heat bath with $d=15$, $d=10$, and $d=5$, respectively.

### v2

@@ Abstract
We consider a three-stroke engine in the microscopic regime, where the working body of the engine is composed of a two-level system. The working body of the engine aims to withdraw heat from the hot heat bath, generate work, and discharge the surplus heat into the cold heat bath through the successive execution of three strokes. In this process, the interaction of the working body with the heat baths is assumed to be energy-conserving and thus can be described by thermal operations. While earlier studies analyzed the optimal performance of this engine when the working body could be transformed by any arbitrary thermal operation, we present closed expressions for the maximum work produced by the engine and the maximum efficiency of the engine when only a restricted class of thermal operations can be implemented on the working body. Furthermore, we explore the engine’s optimal performance under two well-studied classes of restrictions: thermal operations realized via Jaynes-Cummings interaction and thermal operations realizable with finite-sized heat baths. Therefore, on one hand, our results are general, as they reproduce the optimal performance achieved when any arbitrary thermal operation can be implemented on the working body once the restriction is relaxed. On the other hand, our results allow us to determine the engine’s maximum work production and efficiency in a more realistic scenario, where only a restricted class of thermal operations are possible, thereby bringing our findings closer to experimental feasibility.
@@ I Introduction
Within the domain of thermodynamics, heat engines play a pivotal role by converting thermal energy into work. The study of heat engines not only propels technological advancements in energy conversion but also deepens our understanding of nature’s fundamental physical laws. However, the advent of quantum mechanics has ushered in a new era of insight, particularly concerning the foundational underpinnings of thermodynamics on a microscopic scale. This captivating avenue of inquiry traces its origin back to the 1950s, when investigations into the thermodynamic properties of lasers commenced [1, 2, 3]. Since then, a resounding interest has flourished in comprehending the mechanisms of quantum thermal engines as well as quantum thermal machines in general [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], and in formulating microscopic thermodynamic frameworks to characterize their operations [23, 24, 25, 26, 27, 28]. Recently, researchers have realized microscopic heat engines in the lab using a trapped single calcium ion as a working body [29, 30, 31, 32, 33], along with superconducting circuits [34], nitrogen vacancy centers in diamonds [35], atomic [36] and phononic systems [37], and single-electron transistors [38, 39].
@@
The microscopic heat engines can be described as thermodynamic devices that involve small-sized components. In particular, the microscopic heat engines described in the literature involve working bodies that are finite-dimensional and can be as simple as two-level systems [24, 15]. Furthermore, the analysis of the microscopic heat engines has been extended further by considering heat baths of finite size [40, 41, 42, 43, 44]. Based on their way of functioning, the heat engines considered in the microscopic regime can be broadly classified into two classes: continuous engines [11, 7, 45, 46, 5, 47] and stroke-based engines [48, 49, 50, 51]. In a continuous engine, the working body is always in contact with the hot and cold heat baths. As a consequence, the engine withdraws heat from the hot heat bath and converts it into work in a continuous and simultaneous manner. On the other hand, in a stroke-based heat engine, the working body exchanges heat and converts it into work in discrete strokes that are well separated in time. Therefore, in stroke-based engines, the working body repeatedly couples between the hot and cold baths using subsequent strokes.
@@
In this paper, we examine a three stroke heat engine that operates in the microscopic regime introduced in Ref. [52]. By microscopic, we imply that the working body of this engine is a two-level system. The thermodynamic cycle of the engine is composed of three strokes. In the first stroke, the working body aims to withdraw heat from the hot heat bath, followed by the second and third strokes, where the working body converts the heat into work, disposes of the extra amount of heat in the cold heat bath, and completes the cycle. We assume that the interaction between the working body and heat baths is energy-conserving to prevent any unaccounted energy sources from contributing to the work done by the engine. In other words, the engine produces work by harnessing the temperature difference between the two heat baths. This led us to describe the transformation of the state of the working body via a thermal operation [53, 27, 54]. The extraction of work is depicted by a unitary process in which the engine’s working body converts a fraction of the heat extracted from the hot heat bath into work, while the remaining heat is transferred to the cold heat bath.
The three-stroke engine discussed in Ref. [52] presents optimal expressions of the work produced by the engine and its efficiency, assuming that the working body can undergo any arbitrary thermal operation during the first and third strokes. The experimental implementation of any arbitrary thermal operation poses a formidable challenge since the realization of any arbitrary thermal operation assumes infinitely large sized baths with exponentially large amount of degeneracy in the spectrum. Therefore, in a practical scenario, only a restricted class of thermal operations can be realized. This restriction depends on the experimental setup used to realize the three stroke engine. Therefore, achieving the maximum work produced by the engine and the maximum efficiency, as demonstrated in Ref. [52], is very challenging in a realistic scenario.
In this paper, we provide expressions for the maximum work produced by the three-stroke engine and maximum efficiency of it when the engine’s working body undergoes a restricted class of thermal operations. On one hand, our results reproduce the maximum amount of extracted work and the efficiency obtained in Ref. [52] when the restrictions are relaxed. On the other hand, our results allow us to determine the maximum work and efficiency when the thermal operations acting on the working body are restricted to those that can be realized via the Jaynes-Cummings interaction between a two-level working body and a heat bath modeled with a resonant single-mode harmonic oscillator. In addition, we extend our study to a scenario where the restriction on the thermal operations acting on the two-level working body arises due to the presence of a finite-dimensional heat bath modeled with a truncated single-mode harmonic oscillator, rather than an infinitely large heat bath, which is typically required to implement any arbitrary thermal operation.
@@ II.2.1 Heat stroke
This is the initial stroke of the engine, in which the working body interacts with the hot heat bath at thermal equilibrium using energy-conserving unitary $U_{S,H}$, i.e, $[U_{S,H},H_{S}+H_{H}]=0$. Thus, the transformation of the state $\rho_{S}$ is governed by a thermal operation ${\cal E}_{H}$ [27], i.e.,
@@ II.2.2 Work stroke
In the second stroke of the engine, the working body, which resulted from the first stroke, aims to produce work using the heat it drew from the first stroke. The engine extracts work using a cyclic unitary operation that acts on the working body’s state from the first stroke, i.e., ${\cal E}_{H}(\rho_{S})$, and transforms it to the one with lower average energy [8]. A unitary operation is called cyclic if it is generated by a time-dependent Hamiltonian $G(t)$ such that $G(0)=G(\tau)=H_{S}$, where $\tau$ denotes the duration of the work stroke and $H_{S}$ is the Hamiltonian of the working body. We denote the set of such cyclic unitary operations by $\mathcal{V}_{\text{cyc}}$.
In the following, we emphasize an important distinction between our engine and the familiar four-stroke engine operating in Otto cycle (See Fig. 1 (b) of Ref. [55]). In a four-stroke engine operating in an Otto cycle, the thermodynamic cycle consists of two work strokes: the expansion stroke and the compression stroke. After the first work stroke, the free Hamiltonian of the two-level working body changes, and after the second work stroke, it returns to its initial value. In contrast, the proposed three-stroke engine operates with a thermodynamic cycle that includes only one work stroke that extracts work by switching on an interaction that alters the external Hamiltonian of the engine’s working body. The interaction is switched off upon completion of the work stroke. Thus, the Hamiltonian of the working body is reduced to the free Hamiltonian at the end of the work stroke, which is the same as the Hamiltonian at the beginning of the work stroke. In other words, the external driving parameter undergoes cyclic changes, meaning it reverts back to its original value at the end of the work stroke [8, 24].
Therefore, this stroke transforms the working body’s state as follows:
$\mathcal{E}_{H}(\rho_{S})\rightarrow V(\mathcal{E}_{H}(\rho_{S}))V^{\dagger},$ (6)
where $V$ is a cyclic unitary operation. The amount of work produced by the engine is defined as [20, 52]
To maximize the work $W$ over all cyclic unitary operations, we describe the concept of ergotropy introduced in the seminal Ref. [8] by Allahverdyan et al. Consider any arbitrary $d$-dimensional state $\sigma$ having eigendecomposition
$\sigma=\sum_{i=0}^{d-1}\lambda_{i}|\lambda_{i}\rangle\!\langle\lambda_{i}|\quad\text{such that}\quad\forall{i},\;\;\lambda_{i}\geq\lambda_{i+1},$ (8)
associated with Hamiltonian $\mathcal{H}$ with eigendecomposition
$\mathcal{H}=\sum_{j=0}^{d-1}\varepsilon_{j}|\varepsilon_{j}\rangle\!\langle\varepsilon_{j}|\quad\text{such that}\quad\forall{j},\;\;\varepsilon_{j}\leq\varepsilon_{j+1}.$ (9)
Then the ergotropy of the state $\sigma$ defined as follows:
$R(\sigma)$ ::=:= $\max_{\mathcal{U}\in\mathcal{V}_{\text{cyc}}}\text{Tr}\left(\mathcal{H}\left(\sigma-\mathcal{U}\sigma\mathcal{U}^{\dagger}\right)\right)$ (10)
$=$ $\text{Tr}\left(\mathcal{H}\left(\sigma-\mathcal{U}_{\text{max}}\sigma\mathcal{U}_{\text{max}}^{\dagger}\right)\right),$
where $\mathcal{U}_{\text{max}}$ has the following form:
$\mathcal{U}_{\text{max}}=\sum_{j=0}^{d-1}|\varepsilon_{j}\rangle\!\langle\lambda_{j}|.$ (11)
By inspection, we observe that $\mathcal{U}_{\text{max}}$ decreasingly orders the eigenvalues of $\sigma$ such that they correspond to the eigenlevels of Hamiltonian $\mathcal{H}$ arranged in increasing order of energy.
Employing the concept of ergotropy, we can now conclude that when the state of the working body in state ${\cal E}_{H}(\rho_{S})$ transforms via a cyclic unitary operation $V$, the work production by the engine satisfies the following condition:
@@
$\leq$ $\max_{V\in\mathcal{V}_{\text{cyc}}}\text{Tr}\Big{(}H_{S}\big{(}\mathcal{E}_{H}(\rho_{S})-V(\mathcal{E}_{H}(\rho_{S}))V^{\dagger}\big{)}\Big{)}$ (13)
$=$ $R(\mathcal{E}_{H}(\rho_{S})).$ (14)
In other words, the maximum amount of the work produced by the engine is simply equals to the ergotropy of the state of the working body after the first stroke.
@@ II.2.3 Closing the thermodynamic cycle
In the final stroke of the engine, the state of the working body that results from the work stroke $V{\cal E}_{H}(\rho_{S})V^{\dagger}$ interacts with the cold heat bath in thermal equilibrium via an energy conserving unitary operation $U_{S,C}$, i.e., $[U_{S,C},H_{S}+H_{C}]=0$. This transformation corresponds to a thermal operation with respect to the cold heat bath denoted by ${\cal E}_{C}$, i.e.,
@@ II.3 The open cycle engine
Figure 1: The above figure depicts the equivalence between the proposed three-stroke engine and the open-cycle engine introduced in Sec. II.3. On the left panel, we show the proposed three-stroke engine, where the first and third strokes aim to withdraw heat from the hot bath and dispose of the excess heat in the cold bath, whereas the production of work happens during the second stroke. The extracted amount of work can be manifested in terms of the charging of the battery (in this work, we have not considered any explicit battery). The first and third strokes are characterised by thermal operations at the hot and cold inverse temperatures $\beta_{H}$ and $\beta_{C}$, respectively. This proposed engine is equivalent to an open-cycle engine, where the third stroke is characterized by thermalization at some effective inverse temperature $\beta^{v}_{C}$, as shown in the right panel. This equivalence captures the underlying idea behind the proof of theorem 2.
@@ II.4 Mathematical tools
This section provides a brief overview to the mathematical tools associated with thermal operations that are important to evaluate the performance of the engine. A thermal operation can be defined as a complete positive trace preserving (CPTP) map ${\cal E}$, which transforms a system’s state $\rho$ as
@@
where $U$ satisfies the relation $[U,H_{S}+H_{E}]=0$ with $H_{S}$, $H_{E}$ are the Hamiltonians of the system and bath, respectively, and $\gamma_{E}=\frac{e^{-\beta H_{E}}}{\text{Tr}(e^{-\beta H_{E}})}$ is the thermal state of the bath that is thermalized at an inverse temperature $\beta$ [27, 54].
The central task is then described by the following interconversion problem: what is the set of final states that can be achieved from a given state $\rho$ via thermal operations? This set is convex and referred to as the thermal cone of the state $\rho$, denoted by ${\cal T}(\rho)$, which follows straightforwardly from the convexity of thermal operation [56]. Characterizing the necessary and sufficient conditions for the interconversion problem in general is challenging and unknown beyond dimension two [57, 58]. Nonetheless, the seminal work by Horodecki and Oppenheim in Ref. [27] characterizes the necessary and sufficient conditions for the state interconversion problem when $[\rho,H_{S}]=0$. These states are referred to as energy-incoherent states as they do not possess coherence in the energy eigenbasis. For a given energy incoherent state $\rho$, the thermal cone ${\cal T}(\rho)$ is reduced to a polytope, i.e., the set will have a finite number of extremal points, as coherence in the energy eigenbasis can not be generated under thermal operations (See theorem 14 of Ref. [54], [56]). In this case, we refer the set ${\cal T}(\rho)$ as the thermal polytope. The thermal operation that transforms the initial energy incoherent state $\rho$ to an extremal point of the thermal polytope ${\cal T}(\rho)$ is called an extremal thermal operation. Later, we shall see that these extremal points of the set ${\cal T}(\rho)$ and the extremal thermal operations play a crucial role in optimizing the performance of the three stroke engine.
@@
Here, we would like to make an important remark. When we define thermal operations, we suppose that our system of interest is in contact with a very large bath that has an effectively infinite volume. In such an ideal scenario, all possible states within the set ${\cal T}(\rho)$ can be realized. In a more practical context, various restricted subsets of states emerge within the thermal polytope ${\cal T}(\rho)$ as a result of constraints involving the feasibility of thermal operations. In such cases, the range of $\lambda$ in Eq. (25) is more generally, confined between
@@ III.2 Optimal performance of the three stroke engine
The complete proof of this theorem can be found in Appendix C which requires the concept of $\beta$-ordering that we have described in Appendix B [27, 54]. Here, we would like to make an important remark. The expression of the optimal efficiency $\eta_{\textnormal{max}}$ in Eq. (34) reduces to $1+\frac{\lambda_{C}^{\textnormal{max}}}{\lambda_{H}^{\textnormal{max}}\left(1-\lambda_{C}^{\textnormal{max}}\right)}$ when $\omega\rightarrow\infty$, which leads to efficiency more than 1. However, in this regime $W_{\text{max}}$ becomes negative which can be checked by taking the limit $\omega\rightarrow\infty$ in Eq. (LABEL:Opt_performance_W). Note that from Lemma 4 in Appendix A, one can see that in order to produce positive amount of work the efficiency of the engine should be in between 00 and $\eta_{\text{Carnot}}=1-\frac{\beta_{H}}{\beta_{C}}$ i.e.,
$W>0\Rightarrow 0<\eta<\eta_{\text{Carnot}}.$ (35)
By taking negation of Eq. (35), we see the following:
[…]
This implies that the engine fails to produce a positive amount of work when the efficiency of the engine is not in between 00 and $\eta_{\text{Carnot}}$. Therefore, for a given value of $\lambda_{H}^{\textnormal{max}}\in[0,1]$, $\lambda_{C}^{\textnormal{max}}\in[0,1]$, $\beta_{H}$ and $\beta_{C}$ (with the obvious condition $\beta_{C}>\beta_{H}$), we need to suitably choose $\omega$ so that $\eta_{\text{max}}$ lies between 00 and $\eta_{\text{Carnot}}$ to produce a positive amount of work.
Let us look at a scenario where one can perform any thermal operation at inverse temperatures $\beta_{H}$ and $\beta_{C}$ in the first and third strokes, respectively. This implies that one can choose $\lambda^{\text{max}}_{H}=1$ and $\lambda^{\text{max}}_{C}=1$. With these values, one can use the general expressions in Eqs. (LABEL:Opt_performance_W) and (34) to find the expressions for the engine’s maximum work production and maximum efficiency. In that scenario, the optimal work produced by the engine and the optimal efficiency from Eqs. (LABEL:Opt_performance_W) and (34) are reduced to [52]
@@
respectively. Note that, in order to produce a positive amount of work (i.e., $W_{\text{max}}>0$), the necessary and sufficient condition must be
$2>e^{\beta_{H}\omega}+e^{-\beta_{C}\omega}.$ (38)
Therefore, one must follow the protocol in Box 1 to achieve the best work production and efficiency of the stroke-based discrete heat engine with the hot and cold heat baths at inverse temperatures $\beta_{H}$ and $\beta_{C}$, respectively. This specific model of the three-stroke engine has been extensively studied in Ref. [52]. In contrast, our results, which encompass the maximum work production and maximum efficiency given in Eqs. (LABEL:Opt_performance_W) and (34), have broader applicability.
@@
Figure 2: Plot of efficiency ($\eta$) vs $\beta_{C}/\beta_{H}$ with $\omega=0.2/\beta_{H}$. The solid, dashed, dotdashed, and dotted curves represent optimal efficiency, efficiencies for finite heat bath with $d=15$, $d=10$, and $d=5$, respectively.
Figure 3: Plot of $\beta_{H}W$ vs $\beta_{C}/\beta_{H}$ with $\omega=0.2/\beta_{H}$, and $W$ represents the work done by the engine. The solid, dashed, dotdashed, and dotted curves represent optimal work production, work productions for finite heat bath with $d=15$, $d=10$, and $d=5$, respectively.